SPECTROSCOPIC TECHNIQUES FOR COMBUSTION DIAGNOSTICS IN A MICROGRAVITY ENIVI-RONMENT

\underline{NANCY D. PILTCH}, NASA Glenn Research Center, Cleveland OH 44135; JOEL A. SILVER, and DANIEL J. KANE, Southwest Sciences, Inc., Santa Fe NM 87505.

The objective of NASAś microgravity combustion science program is to use that unique environment as a research tool to improve our understanding of fundamental combustion processes and to address spacecraft fire safety. There are multiple challenges in designing instrumentation compatible with this environment including the relatively small allowances for size, weight, power consumption, and time for obtaining data. Additionally experiments must be resistant to vibration and shock loads far in excess of those in terrestrial laboratories; and finally the tests are normally performed with no or minimal operator intervention. Techniques for spectroscopic combustion diagnostics that meet our constraints will be described and some results from drop tower experiments will be presented.